A 2000 kg car, moving along +X direction with a velocity of 72Km/hr, is made to stop by a force

of unknown force-time profile. The car does not recoil and the collision lasts a time-span of à Ž ”t =

0.7s.

[a] Find the initial momentum of the car.

[b] Total impulse felt by the car during the collision.

[c] And average force felt by car during the incident. [Use momentum-impulse theorem].


1
Expert's answer
2021-10-14T11:34:25-0400

[a] The initial momentum is


"p_i=mv_i=40000\text{ kg m/s}."


[b] The total impulse is


"\Delta p=p_i-p_f=p_i-0=40000\text{ kg m/s}."


[c] The force is


"F=\frac{\Delta p}{t}=57.1\text{ N}."